# ERC-SVD: Error-Controlled SVD for Large Language Model Compression

### v3

@@ 1 Introduction
Large language models (LLMs) have emerged as powerful tools, delivering state-of-the-art performance across a wide range of tasks such as text generation, translation, and reasoning. The scaling law [kaplan2020scaling] has driven a trend toward increasingly large models, exemplified by models such as GPT [brown2020language], PaLM [chowdhery2023palm], LLaMA [touvron2023llama], Deepseek [liu2024deepseek], and Qwen [yang2025qwen3], which often contain tens to hundreds of billions of parameters. Despite their powerful capabilities, the enormous scale of LLMs poses serious challenges for efficient deployment due to high computational demands [sheng2023flexgen, zhou2024survey, wang2024model]. This resource burden not only limits deployment on edge devices and consumer-level hardware but also increases the cost and carbon footprint of serving LLMs in production [strubell2020energy, patterson2022carbon].
As the scale of LLMs continues to grow, compression techniques, including weight quantization [frantar2022gptq, lin2024awq, huang2024billm, li2024arb], network pruning [ma2023llm, frantar2023sparsegpt, sunsimple, gao2024disp], knowledge distillation [gu2023minillm, yang2024survey, xu2024survey, zhang2024dual], and low-rank decomposition [hsulanguage, kaushal2023lord, yuan2023asvd, wang2024svd, li2025adasvd], have become increasingly important for the practical deployment of LLMs in resource-constrained environments. The syntactic and semantic correlations acquired during training often induce redundancy in LLM weight matrices, giving rise to a low-rank structure [saha2024compressing]. As a result, singular value decomposition (SVD) provides a principled and effective approach for compressing these matrices with minimal loss. In particular, post-training approaches are gaining traction, as they can significantly reduce memory and compute requirements without the need for expensive retraining, making them especially suitable for scaling up foundation models. Recent studies on post-training SVD-based LLM compression, including ASVD [yuan2023asvd], SVD-LLM [wang2024svd], Basis Sharing [wang2024basis], and AdaSVD [li2025adasvd], have made significant progress in reducing model size while preserving performance, demonstrating the effectiveness of low-rank approximation techniques in compressing LLMs. However, existing methods suffer from two major limitations. First, existing methods ignore the importance of the residual matrix generated during SVD truncation, leading to significant truncation loss. Second, compressing all model layers often results in high layer-wise error and severe error propagation.
@@ 2.1 Techniques for Large Language Model Compression
The growing size of large language models (LLMs) has raised increasing concerns over their computational and memory demands. To address these challenges, a variety of model compression techniques have been proposed. Conventional approaches often require computationally expensive retraining, which is generally impractical due to the substantial computational cost associated with the massive size of LLMs. Consequently, recent efforts have shifted toward more resource-friendly post-training compression techniques [frantar2022gptq, frantar2023sparsegpt, zhu2024survey]. Typically employed approaches include network pruning and weight quantization. And pruning techniques can be classified into unstructured and structured methods. Unstructured pruning removes individual weights based on importance scores. For example, SparseGPT [frantar2023sparsegpt] performs one-shot pruning using second-order approximations without retraining. However, since unstructured pruning retains the original matrix shape, it offers limited inference acceleration and requires specialized hardware. In contrast, structured pruning eliminates entire blocks or channels, enabling compatibility with conventional hardware platforms. LLM-Pruner [ma2023llm] groups dependent linear projections into coupled structures, assigns each group a loss‑aware importance score, prunes the least important groups, and applies LoRA fine‑tuning to restore performance. Additionally, ZipLM [kurtic2023ziplm] prioritizes pruning components that yield the worst trade-off between latency and loss, but this often causes notable performance degradation. Quantization provides another mainstream solution. GPTQ [frantar2022gptq] gradually quantizes and updates each weight using the Hessian matrix to minimize the quantization error. AWQ [lin2024awq] preserves important weight channels by selecting reparameterization coefficients via grid search. Furthermore, BiLLM [huang2024billm] and ARB-LLM [li2024arb] push quantization to the 1-bit level, while still delivering impressive downstream task accuracy. However, these techniques still cause significant performance degradation, especially at low bit-widths, due to the lack of weight and activation adaptation.
@@ 2.2 SVD-based Techniques for Compressing LLMs
Singular value decomposition (SVD) is a commonly employed technique for reducing matrix dimensionality by representing the original matrix as the product of two low-rank factor matrices. Recent research has demonstrated the potential of SVD-based LLM compression methods, yet comprehensive exploration remains limited. Hsu et al. [hsulanguage] propose FWSVD, incorporating Fisher information to weight the importance of parameters. However, it relies on computationally intensive training and was originally applied only to small language models (e.g., BERT [devlin2019bert], ALBERT [lan2019albert]). LoRD [kaushal2023lord] first applies SVD to LLMs by grouping layers to improve efficiency. However, LoRD overlooks the importance of input activations. To address this limitation, Yuan et al. [yuan2023asvd] introduce ASVD, which mitigates the impact of activation outliers by reshaping the weight matrix based on activation distribution, thereby enhancing the precision of the low-rank decomposition. CALDERA [saha2024compressing] exploits the inherent low-rank structure of weight matrices by approximating them through a low-rank, low-precision decomposition. Additionally, SVD-LLM [wang2024svd] establishes a direct mapping between singular values and truncation loss, which means that truncating the smallest singular values leads to minimal truncation loss. Basis Sharing [wang2024basis] achieves more effective LLM compression through sharing parameters across different layers. More recently, AdaSVD [li2025adasvd] proposes an adaptive compensation method, iteratively updating matrices during truncation.
@@ 3 ERC-SVD
The framework of ERC-SVD is illustrated in Figure 2. We first perform SVD to compress the last few layers of the model, while ensuring that the overall compression ratio is satisfied, and compute the corresponding residual matrix. A second SVD is then applied to the residual matrix to obtain its low-rank approximation. The two truncated matrices are subsequently combined to construct the final compressed weight matrices. Compared with other methods [yuan2023asvd, wang2024svd, wang2024basis, li2025adasvd], we utilize the residual matrix during the truncation process to compensate for the truncation loss. In Section 3.1, we first provide a detailed description of how residual compensation effectively works. Subsequently, in Section 3.2, we explain the benefits of partial-layer compression in mitigating error propagation. The pseudocode of ERC-SVD is shown in Algorithm 1, with the pseudocode for residual compensation and partial-layer compression provided in Algorithm 2 and Algorithm 3, respectively.
@@ 3.1 Residual Compensation for SVD Truncation
where $\boldsymbol{U}_{r}\in\mathbb{R}^{m\times r}$ and $\boldsymbol{V}_{r}^{T}\in\mathbb{R}^{r\times n}$ are composed of the top-$r$ left and right singular vectors, respectively, and $\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the corresponding singular values. Following prior post-training SVD-based works [yuan2023asvd, wang2024svd], the optimization objective for SVD truncation in LLMs can be formulated as:
@@
where $\boldsymbol{X}$ denotes the activation of $\boldsymbol{W}$ given an input, and $\mathcal{L}$ is the truncation loss measured by the Frobenius norm. Although previous works [hsulanguage, yuan2023asvd, wang2024svd] have made significant progress in minimizing $\mathcal{L}$, they consistently overlook the residual matrix generated. Based on Equation 3, minimizing the truncation loss reduces to minimizing the discrepancy between the original weight matrix $\boldsymbol{W}$ and its low-rank approximation $\boldsymbol{W}_{r}$. Accordingly, we reformulate the optimization objective as:
@@
Eckart-Young-Mirsky Theorem [horn2012matrix]. Let $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ be a matrix and $r\leq\min(m,n)$ be a given rank. If $\boldsymbol{A}_{r}$ denotes the optimal rank-$r$ approximation of $\boldsymbol{A}$ by SVD, then for any matrix $\boldsymbol{B}$ of rank $r$, the following inequality holds, demonstrating that $\boldsymbol{A}_{r}$ is the optimal rank-$r$ approximation of $\boldsymbol{A}$:
@@ 3.2 Partial-layer Compression for SVD
Previous works [yuan2023asvd, wang2024svd, wang2024basis, li2025adasvd] compress all model layers even if they assign layer-specific ratios based on their relative importance, which often leads to a high layer-wise error, resulting in noticeable performance degradation. We compare the layer-wise error of LLMs across four families with different layer selection strategies; the results of LLaMA-7B and OPT-6.7B are shown in Figure 3, the results of other models are presented in Figure 4. There is a significant error propagation across multiple LLM families, where the error progressively accumulates layer by layer during the forward pass. This phenomenon is particularly pronounced when compressing only the first few layers. A higher layer compression ratio must be applied to these layers, with the constraint of the overall compression ratio. Compressing the model at such an early stage introduces substantial approximation errors, which are then propagated through the remaining layers. Although the later layers are left uncompressed, the forward pass carries the accumulated error through the network. Consequently, in all evaluated LLM families, these early compressed layers exhibit the highest layer-wise error across model layers. In contrast, when only compressing the last few layers, the earlier layers introduce zero error. Although the relatively high compression ratio leads to slightly faster error accumulation, the overall error remains significantly lower than that of compressing all layers.
Interestingly, although the final-layer error converges to a narrow range regardless of how many the last layers are compressed, differences within this range still affect model performance. As illustrated in the left figure of Figure 1, there is a strong correlation between the final-layer error and the average zero-shot accuracy. Motivated by this observation, we select the number of last layers to compress by minimizing the final-layer error, and the layer compression ratio $R_{l}$ must satisfy the following mathematical relation to ensure a fair comparison:
@@
where $N$ is the number of model layers, $R_{o}$ represents the overall compression ratio. As shown in the middle figure of Figure 1, compared to SVD-LLM [wang2024svd], our strategy significantly reduces layer-wise error across all model layers and mitigates error propagation.
@@ 4.1 Setups
We compare our method against five baselines without re-training: conventional SVD, ASVD [yuan2023asvd], SVD-LLM [wang2024svd], Basis Sharing [wang2024basis], and AdaSVD [li2025adasvd]. We evaluate ERC-SVD on ten models spanning five LLM families: LLaMA-2-7B [touvron2023llama2], LLaMA-13B [touvron2023llama], LLaMA-2-13B [touvron2023llama2], LLaMA-30B [touvron2023llama], OPT-6.7B [zhang2022opt], OPT-13B [zhang2022opt], OPT-30B [zhang2022opt], Mistral-7B [jiang2023mistral], Vicuna-7B [chiang2023vicuna], Qwen-3-8B [yang2025qwen3]. For language modeling, we use three benchmark datasets: WikiText-2 [merity2016pointer], PTB [marcus1993building], and C4 [raffel2020exploring]. For zero-shot reasoning and understanding, we evaluate on seven tasks within the LM-Evaluation-Harness framework: OpenbookQA [mihaylov2018can], WinoGrande [sakaguchi2021winogrande], HellaSwag [zellers2019hellaswag], PIQA [bisk2020piqa], MathQA [amini2019mathqa], ARC_e and ARC_c [clark2018think]. Moreover, to demonstrate the generality and effectiveness of ERC-SVD, we further extend our evaluation to vision-language models (VLMs), LLaVA [liu2024improved].
@@ 4.2 Results
Table 1: Overall performance of LLaMA-2-7B [touvron2023llama2] compressed by ERC-SVD and baselines under 20% to 60% compression ratios (“Ratio”), including performance on three language modeling datasets (measured by perplexity ($\downarrow$)) and zero-shot performance on seven common sense reasoning datasets (measured by individual and average accuracy ($\uparrow$)). The best results are marked in bold. NaN denotes evaluation failure due to numerical instability. * refers to results derived from the original paper. - means that results are not available.
@@
We evaluate the performance of LLaMA-2-7B [touvron2023llama2] compressed by ERC-SVD, conventional SVD, and existing post-training baselines under compression ratios ranging from 20% to 60% across ten benchmark datasets. The results for LLaMA-2-7B are presented in Table 1. Across three language modeling datasets, WikiText-2 [merity2016pointer], PTB [marcus1993building], and C4 [raffel2020exploring], ERC-SVD consistently outperforms all baselines across most evaluated compression ratios, with only slight suboptimalities observed in specific cases. In particular, on PTB and C4, the improvements are more pronounced, suggesting that ERC-SVD exhibits stronger generalization capability. In addition, across seven common sense reasoning datasets, ERC-SVD outperforms the strongest existing baseline on most datasets, achieving the highest average accuracy across all evaluated compression ratios and further demonstrating its robustness and overall effectiveness.
@@
To evaluate the generalization ability of ERC-SVD, we apply it to three LLMs from different families, including OPT-6.7B [zhang2022opt], Mistral-7B [jiang2023mistral], and Vicuna-7B [chiang2023vicuna]. As shown in Table 2, under 30% compression ratio, ERC-SVD consistently outperforms all baselines on three language modeling benchmarks across these architectures. Notably, the largest overall improvements are observed on Mistral-7B, with perplexity reductions of 71% on WikiText-2, 45% on PTB, and 46% on C4. We reproduce ASVD and SVD-LLM using their public codebases. While ASVD fails in certain cases due to numerical instability (denoted as NaN in the table), ERC-SVD consistently maintains stable and reliable performance. Additional zero-shot accuracy results comparison about these LLMs are provided in Table 9.
@@
Table 3: Perplexity ($\downarrow$) of larger-scale LLMs under 20% compression ratio on PTB [marcus1993building].

### v4

@@ 1 Introduction
Large language models (LLMs) have emerged as powerful tools, delivering state-of-the-art performance across a wide range of tasks such as text generation, translation, and reasoning. The scaling law [1] has driven a trend toward increasingly large models, exemplified by models such as GPT [2], PaLM [3], LLaMA [4], Deepseek [5], and Qwen [6], which often contain tens to hundreds of billions of parameters. Despite their powerful capabilities, the enormous scale of LLMs poses serious challenges for efficient deployment due to high computational demands [7, 8, 9]. This resource burden not only limits deployment on edge devices and consumer-level hardware but also increases the cost and carbon footprint of serving LLMs in production [10, 11].
As the scale of LLMs continues to grow, compression techniques, including weight quantization [12, 13, 14, 15], network pruning [16, 17, 18, 19, 20], knowledge distillation [21, 22, 23, 24], and low-rank decomposition [25, 26, 27, 28, 29], have become increasingly important for the practical deployment of LLMs in resource-constrained environments. The syntactic and semantic correlations acquired during training often induce redundancy in LLM weight matrices, giving rise to a low-rank structure [30]. As a result, singular value decomposition (SVD) provides a principled and effective approach for compressing these matrices with minimal loss. In particular, post-training approaches are gaining traction, as they can significantly reduce memory and compute requirements without the need for expensive retraining, making them especially suitable for scaling up foundation models. Recent studies on post-training SVD-based LLM compression, including ASVD [27], SVD-LLM [28], Basis Sharing [31], and AdaSVD [29], have made significant progress in reducing model size while preserving performance, demonstrating the effectiveness of low-rank approximation techniques in compressing LLMs. However, existing methods suffer from two major limitations. First, existing methods ignore the importance of the residual matrix generated during SVD truncation, leading to significant truncation loss. Second, compressing all model layers often results in high layer-wise error and severe error propagation.
@@ 2.1 Techniques for Large Language Model Compression
The growing size of large language models (LLMs) has raised increasing concerns over their computational and memory demands. To address these challenges, a variety of model compression techniques have been proposed. Conventional approaches often require computationally expensive retraining, which is generally impractical due to the substantial computational cost associated with the massive size of LLMs. Consequently, recent efforts have shifted toward more resource-friendly post-training compression techniques [12, 17, 32]. Typically employed approaches include network pruning and weight quantization. And pruning techniques can be classified into unstructured and structured methods. Unstructured pruning removes individual weights based on importance scores. For example, SparseGPT [17] performs one-shot pruning using second-order approximations without retraining. However, since unstructured pruning retains the original matrix shape, it offers limited inference acceleration and requires specialized hardware. In contrast, structured pruning eliminates entire blocks or channels, enabling compatibility with conventional hardware platforms. LLM-Pruner [16] groups dependent linear projections into coupled structures, assigns each group a loss‑aware importance score, prunes the least important groups, and applies LoRA fine‑tuning to restore performance. Additionally, ZipLM [33] prioritizes pruning components that yield the worst trade-off between latency and loss, but this often causes notable performance degradation. Quantization provides another mainstream solution. GPTQ [12] gradually quantizes and updates each weight using the Hessian matrix to minimize the quantization error. AWQ [13] preserves important weight channels by selecting reparameterization coefficients via grid search. Furthermore, BiLLM [14] and ARB-LLM [15] push quantization to the 1-bit level, while still delivering impressive downstream task accuracy. However, these techniques still cause significant performance degradation, especially at low bit-widths, due to the lack of weight and activation adaptation.
@@ 2.2 SVD-based Techniques for Compressing LLMs
Singular value decomposition (SVD) is a commonly employed technique for reducing matrix dimensionality by representing the original matrix as the product of two low-rank factor matrices. Recent research has demonstrated the potential of SVD-based LLM compression methods, yet comprehensive exploration remains limited. Hsu et al. [25] propose FWSVD, incorporating Fisher information to weight the importance of parameters. However, it relies on computationally intensive training and was originally applied only to small language models (e.g., BERT [34], ALBERT [35]). LoRD [26] first applies SVD to LLMs by grouping layers to improve efficiency. However, LoRD overlooks the importance of input activations. To address this limitation, Yuan et al. [27] introduce ASVD, which mitigates the impact of activation outliers by reshaping the weight matrix based on activation distribution, thereby enhancing the precision of the low-rank decomposition. CALDERA [30] exploits the inherent low-rank structure of weight matrices by approximating them through a low-rank, low-precision decomposition. Additionally, SVD-LLM [28] establishes a direct mapping between singular values and truncation loss, which means that truncating the smallest singular values leads to minimal truncation loss. Basis Sharing [31] achieves more effective LLM compression through sharing parameters across different layers. More recently, AdaSVD [29] proposes an adaptive compensation method, iteratively updating matrices during truncation.
@@ 3 ERC-SVD
The framework of ERC-SVD is illustrated in Figure 2. We first perform SVD to compress the last few layers of the model, while ensuring that the overall compression ratio is satisfied, and compute the corresponding residual matrix. A second SVD is then applied to the residual matrix to obtain its low-rank approximation. The two truncated matrices are subsequently combined to construct the final compressed weight matrices. Compared with other methods [27, 28, 31, 29], we utilize the residual matrix during the truncation process to compensate for the truncation loss. In Section 3.1, we first provide a detailed description of how residual compensation effectively works. Subsequently, in Section 3.2, we explain the benefits of partial-layer compression in mitigating error propagation. The pseudocode of ERC-SVD is shown in Algorithm 1, with the pseudocode for residual compensation and partial-layer compression provided in Algorithm 2 and Algorithm 3, respectively.
@@ 3.1 Residual Compensation for SVD Truncation
where $\boldsymbol{U}_{r}\in\mathbb{R}^{m\times r}$ and $\boldsymbol{V}_{r}^{T}\in\mathbb{R}^{r\times n}$ are composed of the top-$r$ left and right singular vectors, respectively, and $\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the corresponding singular values. Following prior post-training SVD-based works [27, 28], the optimization objective for SVD truncation in LLMs can be formulated as:
@@
where $\boldsymbol{X}$ denotes the activation of $\boldsymbol{W}$ given an input, and $\mathcal{L}$ is the truncation loss measured by the Frobenius norm. Although previous works [25, 27, 28] have made significant progress in minimizing $\mathcal{L}$, they consistently overlook the residual matrix generated. Based on Equation 3, minimizing the truncation loss reduces to minimizing the discrepancy between the original weight matrix $\boldsymbol{W}$ and its low-rank approximation $\boldsymbol{W}_{r}$. Accordingly, we reformulate the optimization objective as:
@@
Eckart-Young-Mirsky Theorem [36]. Let $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ be a matrix and $r\leq\min(m,n)$ be a given rank. If $\boldsymbol{A}_{r}$ denotes the optimal rank-$r$ approximation of $\boldsymbol{A}$ by SVD, then for any matrix $\boldsymbol{B}$ of rank $r$, the following inequality holds, demonstrating that $\boldsymbol{A}_{r}$ is the optimal rank-$r$ approximation of $\boldsymbol{A}$:
@@ 3.2 Partial-layer Compression for SVD
Previous works [27, 28, 31, 29] compress all model layers even if they assign layer-specific ratios based on their relative importance, which often leads to a high layer-wise error, resulting in noticeable performance degradation. We compare the layer-wise error of LLMs across four families with different layer selection strategies; the results of LLaMA-7B and OPT-6.7B are shown in Figure 3, the results of other models are presented in Figure 4. There is a significant error propagation across multiple LLM families, where the error progressively accumulates layer by layer during the forward pass. This phenomenon is particularly pronounced when compressing only the first few layers. A higher layer compression ratio must be applied to these layers, with the constraint of the overall compression ratio. Compressing the model at such an early stage introduces substantial approximation errors, which are then propagated through the remaining layers. Although the later layers are left uncompressed, the forward pass carries the accumulated error through the network. Consequently, in all evaluated LLM families, these early compressed layers exhibit the highest layer-wise error across model layers. In contrast, when only compressing the last few layers, the earlier layers introduce zero error. Although the relatively high compression ratio leads to slightly faster error accumulation, the overall error remains significantly lower than that of compressing all layers.
Interestingly, although the final-layer error converges to a narrow range regardless of how many the last layers are compressed, differences within this range still affect model performance. As illustrated in the left figure of Figure LABEL:fig:teaser, there is a strong correlation between the final-layer error and the average zero-shot accuracy. Motivated by this observation, we select the number of last layers to compress by minimizing the final-layer error, and the layer compression ratio $R_{l}$ must satisfy the following mathematical relation to ensure a fair comparison:
@@
where $N$ is the number of model layers, $R_{o}$ represents the overall compression ratio. As shown in the middle figure of Figure LABEL:fig:teaser, compared to SVD-LLM [28], our strategy significantly reduces layer-wise error across all model layers and mitigates error propagation.
@@ 4.1 Setups
We compare our method against five baselines without re-training: conventional SVD, ASVD [27], SVD-LLM [28], Basis Sharing [31], and AdaSVD [29]. We evaluate ERC-SVD on ten models spanning five LLM families: LLaMA-2-7B [37], LLaMA-13B [4], LLaMA-2-13B [37], LLaMA-30B [4], OPT-6.7B [38], OPT-13B [38], OPT-30B [38], Mistral-7B [39], Vicuna-7B [40], Qwen-3-8B [6]. For language modeling, we use three benchmark datasets: WikiText-2 [41], PTB [42], and C4 [43]. For zero-shot reasoning and understanding, we evaluate on seven tasks within the LM-Evaluation-Harness framework: OpenbookQA [44], WinoGrande [45], HellaSwag [46], PIQA [47], MathQA [48], ARC_e and ARC_c [49]. Moreover, to demonstrate the generality and effectiveness of ERC-SVD, we further extend our evaluation to vision-language models (VLMs), LLaVA [50].
@@ 4.2 Results
Table 1: Overall performance of LLaMA-2-7B [37] compressed by ERC-SVD and baselines under 20% to 60% compression ratios (“Ratio”), including performance on three language modeling datasets (measured by perplexity ($\downarrow$)) and zero-shot performance on seven common sense reasoning datasets (measured by individual and average accuracy ($\uparrow$)). The best results are marked in bold. NaN denotes evaluation failure due to numerical instability. * refers to results derived from the original paper. - means that results are not available.
@@
We evaluate the performance of LLaMA-2-7B [37] compressed by ERC-SVD, conventional SVD, and existing post-training baselines under compression ratios ranging from 20% to 60% across ten benchmark datasets. The results for LLaMA-2-7B are presented in Table 1. Across three language modeling datasets, WikiText-2 [41], PTB [42], and C4 [43], ERC-SVD consistently outperforms all baselines across most evaluated compression ratios, with only slight suboptimalities observed in specific cases. In particular, on PTB and C4, the improvements are more pronounced, suggesting that ERC-SVD exhibits stronger generalization capability. In addition, across seven common sense reasoning datasets, ERC-SVD outperforms the strongest existing baseline on most datasets, achieving the highest average accuracy across all evaluated compression ratios and further demonstrating its robustness and overall effectiveness.
@@
To evaluate the generalization ability of ERC-SVD, we apply it to three LLMs from different families, including OPT-6.7B [38], Mistral-7B [39], and Vicuna-7B [40]. As shown in Table 2, under 30% compression ratio, ERC-SVD consistently outperforms all baselines on three language modeling benchmarks across these architectures. Notably, the largest overall improvements are observed on Mistral-7B, with perplexity reductions of 71% on WikiText-2, 45% on PTB, and 46% on C4. We reproduce ASVD and SVD-LLM using their public codebases. While ASVD fails in certain cases due to numerical instability (denoted as NaN in the table), ERC-SVD consistently maintains stable and reliable performance. Additional zero-shot accuracy results comparison about these LLMs are provided in Table 9.
@@
Table 3: Perplexity ($\downarrow$) of larger-scale LLMs under 20% compression ratio on PTB [42].